how many atoms of oxygen are present in 1 mole of oxygen?
1
Expert's answer
2013-03-19T10:51:54-0400
To calculate the number of molecules we need to:
For a start let us recall the formula for calculating the number of particles:
n = N/NA

n - the number of moles
N - number of particles
NA - constant Аvogadro (6.02*10^23 mol-1)

From this :

N = NA*n
N =& 6.02*10^23 mol-1 * 1 mol = 6.02*10^23 molecules
Answer:
& Number of molecules = 6.02*10^23 molecules